OpenStudy (anonymous):

medal if you help me POINTS ARE (3,1) AND (-2,-4) A. y + 4 = (x + 2) B. y = -x - 2 C. y - 1 = (x - 3) D. y = x - 2

OpenStudy (shamim):

The formula is y-y1=(y2-y1)/(x2-x1)(x-x1)

OpenStudy (shamim):

x1=3, y1=1

OpenStudy (shamim):

x2=-2, y2=-4

OpenStudy (anonymous):

@xxxghostxxx please stop asking the same questions please just try one by your self its not that challenging just plug in the point to the equation y - y1 = m(x - x1). and (3,1) AND (-2,-4) and x = 3 and -2 and y = 1 and -4

OpenStudy (anonymous):

just plug them in

OpenStudy (anonymous):

@shamim you gave the incorrect equation

OpenStudy (shamim):

no

OpenStudy (shamim):

My formula is like yours

OpenStudy (anonymous):

what is this y-y1=(y2-y1)/(x2-x1)(x-x1) its suppose to be y -y1 = m(x - x1)

OpenStudy (shamim):

Slope m=(y2-y1)/(x2-c1)

OpenStudy (anonymous):

slope is that but you do not divide

OpenStudy (shamim):

Ya

OpenStudy (anonymous):

if that is your equation

OpenStudy (shamim):

C is correct